OpenStudy (anonymous):

Help with multiplying radicals! 6 sqrt {8} times 2 sqrt {4}

OpenStudy (anonymous):

\[6 \sqrt {8} \times 2 \sqrt {4}\]

OpenStudy (saifoo.khan):

48root2

OpenStudy (vijay):

\[48\sqrt{?}\]

OpenStudy (vijay):

\[48\sqrt{?}\]

OpenStudy (saifoo.khan):

\[48\sqrt{2}\]

OpenStudy (vijay):

\[48\sqrt{2}\]

jimthompson5910 (jim_thompson5910):

\[6\sqrt{8}\times2\sqrt{4}\] \[(6\times2)(\sqrt{8}\times\sqrt{4})\] \[12\sqrt{8\times4}\] \[12\sqrt{32}\] \[12\sqrt{16\times2}\] \[12\sqrt{16}\times\sqrt{2}\] \[12\times4\times\sqrt{2}\] \[48\sqrt{2}\] So \[6\sqrt{8}\times2\sqrt{4}=48\sqrt{2}\]

OpenStudy (anonymous):

Please explain how you get the answer! I know it is \[45 \sqrt {2}\] but I don't know how to get that answer

OpenStudy (anonymous):

Thanks for helping!!!

jimthompson5910 (jim_thompson5910):

Answer should be \[48\sqrt{2}\] see steps above on how I got that.

OpenStudy (anonymous):

I got that now thks :D